OpenStudy (anonymous):

Give the x-intercepts of the graph of the equation below. I just need to check my answer

OpenStudy (anonymous):

\[f(x)= \frac{ x ^{2}-5x+4 }{ x ^{2}+5x+6 }\]

OpenStudy (anonymous):

Is it (2,3)?

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

give me a moment i have to go for a few minutes

OpenStudy (anonymous):

okay

OpenStudy (welshfella):

right well you just equate f(x) to zero and solve for x x^2 - 5x + 4 = 0

OpenStudy (welshfella):

the denominator can be ignored because 5x^2 - 5x + 4 = 0 * (x^2 + 5x + 6)

OpenStudy (welshfella):

x^2 - 5x + 4 can be factored

OpenStudy (welshfella):

(2,3) is not correct

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@mustafa2014 can you please help me

OpenStudy (phi):

do you know how to factor the "top" x^2 -5x + 4 ?

OpenStudy (anonymous):

i think so yes

OpenStudy (phi):

list all the pairs of ways to get 4: 1,4 2,2 the + sign on the 4 means the factors will have the same sign the - sign on the -5 means the largest factor is -. that means both factors are - now make both numbers in the pairs negative and add them -1 + -4 = -5 -2 + -2 = -4 if you get a pair that add up to the -5 (from -5x term) those are the factors

OpenStudy (anonymous):

so they're -1, 4

OpenStudy (phi):

so you the factors will be (x-1)(x-4) if we multiply that out using FOIL for example, you get x^2 -5x +4 so that looks good

OpenStudy (anonymous):

yes so the next step would be to factor the bottom?

OpenStudy (phi):

the question asks for the "zeros" the fraction will be zero if the top is zero (and the bottom isn't) all we really want are the ways for the top to be zero.

OpenStudy (phi):

we want (x-1)(x-4) = 0 you can think of that as multiplying two numbers: (x-1) and (x-4) if either is 0, then when we multiply we will get 0 for the top

OpenStudy (phi):

in other words if x-1 = 0 or x-4=0 we will get 0 up top

OpenStudy (anonymous):

i was doing this question using another method. i just put the whole equation equal to 0

OpenStudy (anonymous):

because to find the x intercept shouldn't y be 0?

OpenStudy (phi):

yes, but the "bottom" does not matter if you have a fraction 0/2 or 0/3 or 0/ (pick a number) you get 0 (the only time we care about the bottom is if we get 0/0 ... which is trouble)

OpenStudy (anonymous):

so could i just put the top equal to 0?

OpenStudy (phi):

yes.

OpenStudy (anonymous):

okay so let me figure this out. i need a sec

OpenStudy (anonymous):

i got y=4

OpenStudy (phi):

you want to find the x's that make f(x) = 0 in other words, where the top x^2 -5x + 4 =0 (x-4)(x-1)= 0

OpenStudy (anonymous):

so its (1,4)

OpenStudy (anonymous):

i got it thank you!

OpenStudy (phi):

yes. If you want to be extra careful, be sure the bottom is not zero for x=1 or x=4 (because we don't allow 0/0) but in this case, the bottom will not be 0 for either of those x values. so the zeros of f(x) are at x=1 and x=4